# Estimating LLM Uncertainty with Logits

### v3

@@ 3.4 Considering Logits as Evidence
In this subsection, we present a viable implementation of the LogTokU framework, specifically by modeling AU and EU using a Dirichlet distribution. Inspired by Evidential Deep Learning [22], we treat logits as evidence for each token and model them into a Dirichlet distribution. A naive approach is using the logits of all non-negative values as evidence, while setting those with negative values to no evidence by applying the ReLU activation function. However, unlike conventional classification networks, the number of candidates generated by an LLM (that is, tokenizer size) is significantly large, with a considerable proportion of tokens having extremely low logits, and these tokens should be discarded as noise [28]. Therefore, this paper focuses on the distribution of main candidates with high logits. Specifically, we sample the top $K$ tokens with the largest logits to model a Dirichlet distribution:
@@
3.5 Why Logits can Capture EU
Why evidential modeling can capture EU. Due to there is no ground-truth of the EU, uncertainty-aware model training (such as evidential deep learning [22]) does not obtain the EU through direct alignment. Instead, it is typically achieved by assuming that all scenarios should intuitively have a high EU for the model, and the EU is gradually eliminated during the model’s learning process. Therefore, there is a regularization constraint on the growth of the evidences (logits). This process consists of two parts: classification loss and regularization for constraining smaller logits. In other words, due to the trade-off between correct classification and logits regularization, logits are only allowed to be larger when the model is highly confident in its predictions, thereby characterizing the magnitude of the EU. Taking the training loss of typical evidential modeling ($\mathcal{L}_{\text{EVD}}$) as an example:
$\mathcal{L}_{\text{EVD}}=\underbrace{-\left(\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{z_{j}+1}{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}\right)}_{\text{Evidential deep learning classification loss}}\underbrace{+\mathcal{L}_{\text{reg}}}_{\text{logits constrainer}},$ (6)
where $z_{j}$ is the predictive logit of class $j$ and $\mathcal{L}_{\text{reg}}$ indicates the regularization for constraining the logits scale.
Cross-entropy can be interpreted as evidential modeling. Intuitively, the cross-entropy loss is not affected by the scale of logits; adding a constant to all elements of the logits vector does not change the probability values. For instance, both [20, 18] and [30, 28] are mapped to the probability distribution [0.88, 0.12] after applying the softmax function. However, intuition is not always trustworthy. During the actual gradient descent training of models, the scale of logits is constrained. In fact, as shown in Eq. 7, the cross-entropy loss can be decomposed into a format similar to the evidential deep learning loss function, which naturally penalizes shifting logits scales during gradient updates.
$\mathcal{L}_{\text{LLM}}$ $=-\sum_{j}^{|\bm{Y}|}y_{j}\log p_{j}$ (7)
$=-\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{e^{z_{j}}}{\sum_{j^{\prime}}^{|\bm{Y}|}e^{z_{j^{\prime}}}}$
$=-\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{e^{(z_{j}+1})}{\sum_{j^{\prime}}^{|\bm{Y}|}e^{(z_{j^{\prime}}+1})}$
$=\underbrace{-\left(\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{z_{j}+1}{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}\right)}_{\text{Evidential deeplearning classification loss}}$
@@ 3.5 Why Logits can Capture EU
$\quad\underbrace{-\log\frac{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}{\sum_{j^{\prime}}^{|\bm{Y}|}e^{(z_{j^{\prime}}+1)}}}_{\text{logits constrainer}},$
In other words, absolute logits scale does not change probability distribution, but it affects optimization, which is the theoretical reason why logits can be considered as evidence similar to uncertainty-aware modeling.
@@ 5.1 LogTokU-guided Response Uncertainty Estimation
$\mathcal{R}(a_{t})=\frac{1}{\text{AU}(a_{t})\cdot\text{EU}(a_{t})},$ (9)
@@ 5.2.2 Results Analyses
6 Theoretical Analysis
[…]
where $L^{\tau^{a}}$ is the loss on the sample with the next token label $\tau^{a}$. Consider cases where multiple distinct answers to the same question appear in the training set, the situation becomes different. For example, $\tau^{b}$ (“George”) is the label in another sample with the same question. When the model is simultaneously fine-tuned on both samples, the gradient update for the model will be:
$\quad=\underbrace{-y_{a}^{\tau^{a}}\frac{1}{\Omega_{a}^{\tau^{a}}}\nabla_{\mathcal{M}}\Omega_{a}^{\tau^{a}}-y_{b}^{\tau^{b}}\frac{1}{\Omega_{b}^{\tau^{b}}}\nabla_{\mathcal{M}}\Omega_{b}^{\tau^{b}}}_{\text{(1) maximizing the probability of annotated answer}}$
where $\Omega_{a}^{\tau^{a}}=\frac{\exp(\mathcal{M}(\tau^{a}|\bm{q},\bm{a}_{t-1}))}{\sum_{m=1}^{|\bm{Y}|}\exp(\mathcal{M}(\tau^{m}|\bm{q},\bm{a}_{t-1}))}$, and $y_{a}^{\tau^{m}}$ indicates the next token label of a training sample with ground-truth label ${\tau^{a}}$, that is, we have $y_{a}^{\tau^{a}}=1$ and $y_{a}^{\tau^{b}}=0$. In particular, when $\mathcal{M}$ is in a certain state during training, we have $\Omega_{a}^{\tau^{a}}=\Omega_{b}^{\tau^{a}}$, and we make distinctions to facilitate the reader’s understanding here. As we can see, for scenarios with multiple answers, the training objective can be divided into three parts: (1) For each sample, increase the probability of its own annotation in the output distribution. (2) For each sample, decrease the probability of another sample’s annotation in the output distribution. Note: This part leads to the issue where probability cannot anymore capture the reliability of LLM responses, as different correct answers tend to reduce the probability of other correct answers, making low probabilities cannot indicates low reliability. (3) For both samples, decrease the probability of other outputs not present in the annotations in the output distribution.
@@ Appendix F Limitations
Appendix G Theoretical Proof of Why Logits can Capture Epistemic Uncertainty

### v4

@@ 3.4 Considering Logits as Evidence
In this subsection, we present a viable implementation of the LogTokU framework, specifically by modeling AU and EU using a Dirichlet distribution. Inspired by Evidential Deep Learning [22], we treat logits as evidence for each token and model them into a Dirichlet distribution. A naive approach is using the logits of all non-negative values as evidence, while setting those with negative values to no evidence by applying the ReLU activation function. However, unlike conventional classification networks, the number of candidates generated by an LLM (that is, tokenizer size) is significantly large, with a considerable proportion of tokens having extremely low logits, and these tokens should be discarded as noise [28]. Therefore, this paper focuses on the distribution of main candidates with high logits. Specifically, we select the logits of tokens with the top $K$ largest logits to model a Dirichlet distribution:
@@
3.5 Theoretical Analysis: Why Logits can Capture Uncertainty
Intuitively, the cross-entropy loss is not affected by the scale of logits; adding a constant to all elements of the logits vector does not change the probability values. For instance, both [20, 18] and [30, 28] are mapped to the probability distribution [0.88, 0.12] after applying the softmax function. However, intuition is not always trustworthy. During the actual gradient descent training of models, the scale of logits is constrained. The cross-entropy loss can be decomposed into a format similar to the evidential deep learning loss function, which naturally penalizes shifting logits scales during gradient updates:
$\mathcal{L}_{\text{LLM}}$ $=\underbrace{-\left(\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{z_{j}+1}{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}\right)}_{\text{Evidential deeplearning classification term}}$ (6)
@@ 3.5 Theoretical Analysis: Why Logits can Capture Uncertainty
$\underbrace{-\log\frac{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}{\sum_{j^{\prime}}^{|\bm{Y}|}e^{(z_{j^{\prime}}+1)}}}_{\text{Evidence regularization}},$
which is very close to the format of training loss in evidential modeling shown in the Appendix B.3, and the following analysis show the detailed differences and corresponding impacts between evidential modeling loss and CE loss.
Bridging the positivity gap. A critical distinction lies in evidence parameterization: evidential models enforce $z_{j}\geq 0$ through architectural constraints (e.g., ReLU activations), whereas standard CE allows unbounded logits. However, the exponential transformation in softmax ($e^{z_{j}}$) renders negative logits equivalent to weak evidence (values near 0), while positive logits represent strong evidence. Thus, both frameworks ultimately operate in a non-negative evidence space, with CE’s logits implicitly encoding $\log(\text{evidence}+\epsilon)$ where $\epsilon\to 0$. This logarithmic relationship preserves the uncertainty-quantification mechanism while relaxing explicit positivity constraints.
The role of gradient dynamics. The purported scale invariance of CE holds only for completed training – during optimization, gradient updates actively regulate logit magnitudes. For misclassified samples, the $\log\sum e^{z_{j}}$ term exerts downward pressure on all logits, while the $-y_{j}z_{j}$ term selectively boosts the true class logit. This creates an implicit evidence threshold: logits only grow when the model can simultaneously increase the true class evidence while suppressing competing terms, mirroring the explicit trade-off in evidential learning.
@@ 5.1 LogTokU-guided Response Uncertainty Estimation
$\mathcal{R}(a_{t})=-{\text{AU}(a_{t})\cdot\text{EU}(a_{t})},$ (8)
@@ B.2 Max Token Probability Can Not Represent Risk of A Wrong Answer
B.3 How can Evidential Modeling Capture Uncertainty
Due to the absence of ground-truth uncertainty, uncertainty-aware training methods like evidential deep learning [22] do not directly supervise EU. Instead, they implicitly learn to quantify uncertainty by imposing inductive biases during optimization. This is achieved through a dual mechanism: (1) a classification term that aligns predictions with labels, and (2) a regularization term that constrains evidence accumulation. Specifically, the model is penalized for accumulating excessive evidence (large logits) unless it achieves high confidence in predictions. The training loss $\mathcal{L}_{\text{EVD}}$ exemplifies this principle:
$\mathcal{L}_{\text{EVD}}=\underbrace{-\left(\sum_{j}^{|\bm{Y}|}y_{j}\log\frac{z_{j}+1}{\sum_{j^{\prime}}^{|\bm{Y}|}(z_{j^{\prime}}+1)}\right)}_{\text{Classification term}}\underbrace{+\mathcal{L}_{\text{reg}}}_{\text{Evidence regularization}},$ (15)
where $z_{j}$ represents non-negative evidence parameters for class $j$, and $\mathcal{L}_{\text{reg}}$ penalizes premature evidence accumulation.
@@ Appendix F Limitations
Appendix G Theoretical Analyze of Why Logits can Capture Epistemic Uncertainty